# Rethinking Early-Fusion Strategies for Improved Multispectral Object Detection

### v1

@@ III-D Loss Function
| mAP | 49.10 | 52.40 | 53.30 | 53.00 | 52.20 | 52.70 | 53.40 | 53.10 | 53.50 | 53.10 | 52.50 | 53.40 | 54.20 |
| mAP50 | 77.30 | 81.90 | 81.90 | 80.90 | 81.60 | 81.50 | 81.70 | 82.00 | 81.70 | 82.20 | 81.00 | 81.90 | 83.40 |
| Person | 79.30 | 68.40 | 76.70 | 76.30 | 73.60 | 74.50 | 76.50 | 76.70 | 75.30 | 77.00 | 79.10 | 77.00 | 79.90 |
| Car | 87.90 | 90.80 | 91.00 | 91.00 | 90.70 | 91.10 | 91.40 | 90.80 | 91.00 | 91.10 | 90.50 | 91.20 | 91.80 |
| Bus | 87.20 | 92.20 | 90.00 | 90.10 | 90.70 | 91.60 | 89.40 | 90.10 | 89.40 | 91.20 | 89.40 | 90.70 | 89.20 |
| Motor | 70.00 | 74.00 | 72.60 | 69.20 | 74.80 | 73.50 | 72.80 | 73.30 | 73.30 | 72.20 | 70.30 | 71.30 | 76.20 |
| TrafficLight | 55.90 | 80.30 | 77.40 | 75.40 | 76.90 | 74.80 | 77.20 | 78.20 | 77.40 | 77.60 | 72.70 | 77.70 | 78.30 |
| Truck | 83.40 | 85.70 | 83.70 | 83.10 | 82.90 | 83.40 | 82.90 | 82.90 | 83.90 | 84.10 | 84.00 | 83.60 | 85.30 |
@@
| mAP | 34.90 | 36.10 | 38.30 | 38.60 | 37.10 | 37.10 | 38.90 | 38.00 | 38.20 | 38.90 | 39.10 | 38.70 | 41.50 |
| mAP50 | 57.20 | 60.20 | 62.00 | 61.90 | 61.00 | 60.80 | 62.40 | 61.10 | 61.30 | 62.20 | 61.90 | 61.90 | 66.80 |
| Person | 74.60 | 55.90 | 72.20 | 71.90 | 67.30 | 67.60 | 72.30 | 71.70 | 70.50 | 72.50 | 75.50 | 72.40 | 77.60 |
| Car | 80.20 | 84.80 | 85.50 | 85.60 | 84.90 | 85.20 | 85.90 | 85.50 | 85.80 | 85.50 | 85.00 | 85.50 | 87.40 |
| Bus | 58.30 | 65.70 | 58.60 | 61.80 | 61.60 | 59.80 | 61.40 | 58.30 | 61.30 | 61.50 | 60.90 | 60.10 | 65.10 |
| Motor | 48.00 | 45.10 | 49.10 | 47.60 | 49.00 | 48.70 | 49.60 | 45.80 | 44.60 | 47.50 | 46.80 | 50.80 | 55.60 |
| TrafficLight | 27.30 | 56.80 | 49.80 | 48.70 | 49.10 | 51.20 | 48.60 | 50.90 | 49.70 | 50.80 | 46.90 | 48.00 | 54.90 |
| Truck | 54.80 | 52.70 | 56.70 | 55.50 | 53.80 | 52.60 | 56.80 | 54.70 | 55.80 | 55.70 | 56.70 | 54.70 | 60.30 |
@@
| mAP50 | 67.20 | 61.40 | 61.54 | 69.62 | 72.39 | 72.90 | 73.20 | 73.58 | 74.60 | 76.20 | 79.20 | 80.00 | 83.76 | 84.80 |
| Bicycle | 60.50 | 50.26 | 49.43 | 55.53 | 55.77 | - | 57.40 | 56.20 | 60.04 | - | - | 67.70 | 73.49 | 79.80 |
| Car | 83.60 | 70.63 | 70.72 | 82.33 | 84.91 | - | 86.50 | 85.80 | 85.52 | - | - | 89.00 | 90.14 | 92.80 |
| Person | 57.60 | 63.31 | 64.47 | 71.01 | 74.49 | - | 75.60 | 78.74 | 78.24 | - | - | 83.20 | 87.65 | 81.20 |
IV Experiments
IV-A Experimental Setup
Datasets. Our experiments are conducted on the M3FD dataset [1] and FLIR dataset [35]. M3FD dataset contains 4200 pairs of RGB and thermal images. These image pairs are well aligned. The dataset contains 6 classes of objects: ‘Person’, ‘Car’, ‘Bus’, ‘Motorcycle’, ‘Traffic Light’, and ‘Truck’. Since this dataset doesn’t provide unified data splits, previous works have used a random splitting approach to determine the train and validation sets [1]. However, images in this dataset are sampled from video sequences, meaning that two adjacent frames may contain identical content. In this context, the random splitting approach results in information leakage between the train and validation sets. To address this problem, we first manually divide the dataset into 73 video segments based on different scenes. Then, we collect the first 70% of images in each video segment as the train set and the remaining images as the validation set. Finally, we obtain 2905 and 1295 pairs of RGB-T images in the train and validation sets, respectively. We name this data split ‘M3FD-zxSplit’ and release it to the public111https://github.com/XueZ-phd/Efficient-RGB-T-Early-Fusion-Detection. For the performance evaluation in Section IV-B, we use this data split. When comparing with state-of-the-art approaches in Section IV-C, we employ both ‘M3FD-zxSplit’ and random splitting. Our random splitting refers to randomly selecting 80% images as the train set and the remaining images as the validation set. FLIR dataset originally contains unaligned RGB-T image pairs. The work [57] develops a data-processing approach to align these images and obtain 7381, 1056, and 2111 image pairs in the train, validation, and test sets. This dataset contains 3 classes: ‘Person’, ‘Bicycle’ and ‘Car’.
Implementation Details. We incorporate our three key modules into commonly-used one-stage detectors, including RetinaNet [30], GFL [31], and YOLOv5 [2]. For RetinaNet and GFL, we adopt the implementations in MMDetection toolbox [58]. For YOLOv5, we use its official implemtation [2]. We keep the training setting consistent with the corresponding baselines.
IV-B Performance Evaluation
Table I and Table II present the performance of our method on the M3FD [1] and FLIR [35] datasets. Key observations include: (1) the medium-fusion strategy adds more parameters and FLOPs compared to the early-fusion strategy; (2) the medium-fusion strategy achieves better performance compared to single-modality inputs, whereas the plain early-fusion strategy does not consistently improve performance; (3) our EME method, incorporating the ShaPE module, weakly supervised learning, and CoreKD techniques into the plain early-fusion strategy, achieves significant performance improvement without significantly increasing parameters and FLOPs; (4) the inference time of our EME method is longer than that of the baseline method, since the structural similarity computation process has not been optimized when calculating the self-gating mask; (5) our EME method can outperform the medium-fusion strategy in both performance and efficiency to some extent.
IV-C Comparison with the State-of-The-Art Approaches
@@
We use the one-stage YOLOv5 [2] detector as the baseline, and incorporate our proposed modules to construct the effective multispectral early-fusion (EME) model. Table III and Table IV compares our EME and previous state-of-the-art approaches on M3FD [1] and FLIR [35] datasets.
@@
In Table III, we compare our EME with 10 state-of-the-art image-fusion-based object detection approaches [40, 24, 41, 25, 42, 43, 44, 45, 1, 46]. We first generate fused images based on their official implementations, and then train YOLOv5 [2] using these fused images with the same training settings. The results show that our EME achieves state-of-the-art performance. We observe that the results in Table III (a) are obviously better than those in Table III (b). This demonstrates that random splitting causes information leakage and makes it difficult to improve performance. Fig. 13 presents an example scene for visualization.
@@
In Table IV, we compare our EME with 13 multispectral object detection approaches. These approaches include (1) medium-fusion strategies, such as CBF [47], MCG [48], MUN [48], CFR [50], GAFF [15], SMPD [52], MSAT [54], CSAA [55], and MFPT [56]; (2) domain adaptation and single-modality detection approaches, such as ODS [49], BU [51], and ThDe [53]; and (3) late-fusion strategy [22]. The results show that our EME also achieves state-of-the-art performance on the FLIR dataset [35].
@@ V Conclusions
In this paper, we have proposed the effective multispectral early-fusion (EME) detector, which achieves both high performance and efficiency. We identify and address performance obstacles such as information interference, domain gap, and weak feature presentation, proposing solutions including shape-priority early-fusion modules, weakly supervised learning, and core knowledge distillation. Extensive experiments demonstrate the effectiveness and efficiency of our EME.

### v2

@@ IV-C Performance Evaluation of Variants
mAP | 49.10 | 52.40 | 53.30 | 53.00 | 52.20 | 52.70 | 53.40 | 53.10 | 53.50 | 53.10 | 52.50 | 53.40 | 54.00±0.28 |
mAP50 | 77.30 | 81.90 | 81.90 | 80.90 | 81.60 | 81.50 | 81.70 | 82.00 | 81.70 | 82.20 | 81.00 | 81.90 | 82.90±0.37 |
Person | 79.30 | 68.40 | 76.70 | 76.30 | 73.60 | 74.50 | 76.50 | 76.70 | 75.30 | 77.00 | 79.10 | 77.00 | 79.53±0.26 |
Car | 87.90 | 90.80 | 91.00 | 91.00 | 90.70 | 91.10 | 91.40 | 90.80 | 91.00 | 91.10 | 90.50 | 91.20 | 91.90±0.29 |
Bus | 87.20 | 92.20 | 90.00 | 90.10 | 90.70 | 91.60 | 89.40 | 90.10 | 89.40 | 91.20 | 89.40 | 90.70 | 89.80±0.45 |
Motor | 70.00 | 74.00 | 72.60 | 69.20 | 74.80 | 73.50 | 72.80 | 73.30 | 73.30 | 72.20 | 70.30 | 71.30 | 74.87±0.95 |
TrafficLight | 55.90 | 80.30 | 77.40 | 75.40 | 76.90 | 74.80 | 77.20 | 78.20 | 77.40 | 77.60 | 72.70 | 77.70 | 77.40±1.13 |
Truck | 83.40 | 85.70 | 83.70 | 83.10 | 82.90 | 83.40 | 82.90 | 82.90 | 83.90 | 84.10 | 84.00 | 83.60 | 84.00±0.93 |
@@
mAP | 34.90 | 36.10 | 38.30 | 38.60 | 37.10 | 37.10 | 38.90 | 38.00 | 38.20 | 38.90 | 39.10 | 38.70 | 41.10±0.29 |
mAP50 | 57.20 | 60.20 | 62.00 | 61.90 | 61.00 | 60.80 | 62.40 | 61.10 | 61.30 | 62.20 | 61.90 | 61.90 | 66.23±0.40 |
Person | 74.60 | 55.90 | 72.20 | 71.90 | 67.30 | 67.60 | 72.30 | 71.70 | 70.50 | 72.50 | 75.50 | 72.40 | 77.23±0.26 |
Car | 80.20 | 84.80 | 85.50 | 85.60 | 84.90 | 85.20 | 85.90 | 85.50 | 85.80 | 85.50 | 85.00 | 85.50 | 87.13±0.19 |
Bus | 58.30 | 65.70 | 58.60 | 61.80 | 61.60 | 59.80 | 61.40 | 58.30 | 61.30 | 61.50 | 60.90 | 60.10 | 62.33±1.96 |
Motor | 48.00 | 45.10 | 49.10 | 47.60 | 49.00 | 48.70 | 49.60 | 45.80 | 44.60 | 47.50 | 46.80 | 50.80 | 55.33±0.21 |
TrafficLight | 27.30 | 56.80 | 49.80 | 48.70 | 49.10 | 51.20 | 48.60 | 50.90 | 49.70 | 50.80 | 46.90 | 48.00 | 55.33±0.37 |
Truck | 54.80 | 52.70 | 56.70 | 55.50 | 53.80 | 52.60 | 56.80 | 54.70 | 55.80 | 55.70 | 56.70 | 54.70 | 60.10±0.16 |
@@
mAP50 | 67.20 | 61.40 | 61.54 | 69.62 | 72.39 | 72.90 | 73.20 | 73.58 | 74.60 | 76.20 | 79.20 | 80.00 | 83.76 | 84.63±0.12 |
Bicycle | 60.50 | 50.26 | 49.43 | 55.53 | 55.77 | - | 57.40 | 56.20 | 60.04 | - | - | 67.70 | 73.49 | 79.73±0.09 |
Car | 83.60 | 70.63 | 70.72 | 82.33 | 84.91 | - | 86.50 | 85.80 | 85.52 | - | - | 89.00 | 90.14 | 92.67±0.12 |
Person | 57.60 | 63.31 | 64.47 | 71.01 | 74.49 | - | 75.60 | 78.74 | 78.24 | - | - | 83.20 | 87.65 | 81.27±0.09 |
IV-D Comparison with the State-of-The-Art Approaches
@@
We use the one-stage YOLOv5 [2] detector as the baseline, and incorporate our proposed modules to construct the effective multispectral early-fusion (EME) model. Table VI and Table VII compares our EME and previous state-of-the-art approaches on M3FD [1] and FLIR [39] datasets.
@@
In Table VI, we compare our EME with 10 state-of-the-art image-fusion-based object detection approaches [51, 30, 52, 31, 53, 54, 55, 56, 1, 57]. We first generate fused images based on their official implementations, and then train YOLOv5 [2] using these fused images with the same training settings. The results show that our EME achieves state-of-the-art performance. We observe that the results in Table VI (a) are obviously better than those in Table VI (b). This demonstrates that random splitting causes information leakage and makes it difficult to improve performance. Fig. 14 presents an example scene for visualization. Compared to other approaches, a weakness of our EME detector is that it doesn’t generate a fused image for direct visualization. This is because our method focuses on detection rather than image fusion. We will address this issue in future work.
@@
In Table VII, we compare our EME with 13 multispectral object detection approaches. These approaches include (1) medium-fusion strategies, such as CBF [58], MCG [59], MUN [59], CFR [61], GAFF [15], SMPD [63], MSAT [65], CSAA [66], and MFPT [67]; (2) domain adaptation and single-modality detection approaches, such as ODS [60], BU [62], and ThDe [64]; and (3) late-fusion strategy [28]. The results show that our EME also achieves state-of-the-art performance on the FLIR dataset [39].
IV-E Comparison of Inference Efficiency
We compare the inference efficiency of our EME method with previous state-of-the-art approaches on an edge device: the NVIDIA AGX Orin with 64GB of memory. We select open-source approaches for comparison and adopt YOLOv5-small as the baseline detector for all methods.
Table VIII presents the FLOPs, number of parameters, and inference time for various methods. Experimental results show that our EME method is the fastest. Interestingly, we notice that a reduction in FLOPs does not directly lead to a similar reduction in the inference time of an approach. This phenomenon may be attributed to the frequent memory access by operators, as confirmed in PConv[68]. This observation inspires us to further speed up our EME method by reducing memory access in the future.
@@ V Conclusions
In this paper, we propose the effective multispectral early-fusion (EME) detector, which achieves both high performance and efficiency. We identify and address performance obstacles in a plain early-fusion strategy, such as information interference, domain gaps, and weak feature representation, by proposing solutions including shape-priority early-fusion modules, weakly supervised learning, and core knowledge distillation. Extensive experiments on representative datasets demonstrate the effectiveness and efficiency of our EME detector.
The main advantage of our EME detector is that it improves the performance of an efficient single-branch early-fusion strategy without significantly increasing its computational burden. We demonstrate that our EME detector has similar FLOPs and parameters to a plain early-fusion strategy, while achieving better performance than a cumbersome two-branch structure. We also show that our EME detector has higher inference efficiency than the two-branch structure on an edge device: the NVIDIA AGX Orin.
A limitation of our current EME detector is its inefficient two-stage training paradigm in the knowledge distillation technique. In the future, we will work towards an optimized one-stage paradigm to accelerate the training process and further improve detection accuracy.